OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = the quantity of 3x-1/2 solve for g(f(3)). A. 5 B. 7 C. 9 D. 10

OpenStudy (jhannybean):

clarification: \(g(x) = \frac{3x-1}{2}\) or is it \(g(x) = 3x-\frac{1}{2}\) ?

OpenStudy (anonymous):

the first one

OpenStudy (jhannybean):

Ok :)

OpenStudy (jhannybean):

We can find what f(3) is first

OpenStudy (jhannybean):

f(x) = 2x + 1 \(\implies\) f(3) = 2(3) + 1

OpenStudy (jhannybean):

@mariyakelly

OpenStudy (anonymous):

okay I don't know how to solve that @Jhannybean

OpenStudy (jhannybean):

What is 2 * 3?

OpenStudy (jhannybean):

then add 1.